INFRARED EMISSION STUDIES OF THE A^{3}Σ^{−}-X^{3}Π ELECTRONIC TRANSITION OF SiC

\underline{M. N. DEO} and K. KAWAGUCHI, Nobeyama Radio Observatory, Minamimaki, Minamisaku, Nagano, 384-1305, Japan.

The gas phase infrared emission spectrum of the A^{3}Σ^{−} - X^{3}Π electronic transition of SiC has been observed using Fourier transform spectrometer. The SiC radical was generated by a dc discharge in a flowing mixture of (CH_{3})_{6}Si_{2} and He. Three bands 1-0 (4577.8 cm^{−1}), 0-0 (3723.1 cm^{−1}) and 0-1 (2769.8 cm^{−1}) have been observed, out of these the last two were observed for the first time^{a}. Altogether more than 1100 transitions have been assigned and these data were simultaneously least squares fitted and obtained the molecular constants for SiC in the A^{3}Σ^{−} and X^{3}Π electronic states. The vibrational frequency ν_{0} in the A state was determined to be 854.6994(7) cm^{−1}, which is close to the matrix result^{b} 854.2 cm^{−1}.

a4577.8 cm^{−1} band was first observed by Brazier et.al.[J. Chem. Phys., 91, 7384(1989)] and they identified it as 0-0 band.

bM. Grutter, P. Freivogel and J. P Maier, J. Phys. Chem. A, 101, 275(1997)